?? Ar- duino ?? ? train.eps, width =

100mmP owerT rainmovementflowschemebetweenparts. (F_{eng}) (F_{drag}) (F_{g}) (F_{rr}) (F_{b})

?? [ columnsep =

10mm][plane](0, −1)coordinate(base) − −coordinate[pos = 0.5](mid) + +(30 : 3)coordinate(top)| − (base) − −cycle; α +y +x α N F_{b} F_{rr} F_{drag}

F_{eng} F_{g}

F_{r} = F_{eng}+F_{drag}+F_{g}+F_{rr}+F_{b} (1)

?

F_{drag} = 12×ρ×C_{d}×A×v^{2}

(2) ρ ⟶

[kg/m^{3}]

C_{d} ⟶ A ⟶ [m^{2}] v ⟶ [m/s]

F_{rr} = C_{rr}×N (3)

C_{rr} ⟶ N ⟶ [N]

N = m×g×cos(θ)

(4) m ⟶ [kg] g ⟶

[m/s^{2}]

θ ⟶ ?

C_{rr} = 0.005+1P ×(0.01+0.0095×(v100)^{2})

(5)

P ⟶ [bar] v ⟶ [km/h]

F_{b} = percent×F_{b_{max}} (6)

percent ⟶ F_{b_{max}} ⟶ [N]

F_{g} = −m×g×sin(θ)

(7) m ⟶ [kg] g ⟶

[m/s^{2}]

θ ⟶ ? rpm_{c}urve.eps, width =

100mmT orquevsRP Mofa1999DodgeNeonDOHCengine. ?? ?

τ = throttlepercent×τ_{max}@RP M (8)

throttlepercent ⟶ τ_{max}@RP M ⟶ [N· m]

F_{eng} = τ × x_{g} × x_{d} × η_{trans}R_{tire}

(9) τ ⟶ [N·

m] x_{g} ⟶ x_{d} ⟶ η_{trans} ⟶

R_{tire} ⟶ [m]

?? gui_{s}imulator.eps, width =

80mmP arametersGraphicalUserInterfaceofthesimulator. Physical Characteristics Car Characteristics